Two trains are traveling on the same track and in the same direction. The first train, which is behind the second train, blows a horn whose frequency is 266 Hz. The second train detects a frequency of 284 Hz. If the speed of the second train is 13.8 m/s, what is the speed of the first train?
Answer #1

- 2 f = fx (sound - train-2 g (Dopplers Effect, - I sound - train ) 284 = 266* 343 -13.8 7 L 343 - VI = v = 34-66 m/s Answe